How many milliliters of dry CO2, measured at STP, could be evolved in the reaction between 17.0 mL of 0.293 M NaHCO3 and 31.0 mL of 0.864 M HCl?


1
Expert's answer
2021-10-24T01:45:03-0400

First find out the mmoles of NaHCO3and HCl

moles = Molarity × volume

moles of NaHCO3= ( 0.270M ×16.2mL ) = 4.374 moles

moles of HCl = ( 0.753M × 34.2mL ) = 25.75 moles

According to the chemical reaction equation,

1mol of NaHCO3 reacts with 1mol of HCl to Produces 1mol of CO2

Then,

4.374mol of NaHCO3 react with 4.374mol of HCl to Produces 4.374mol of CO2

Calculate the volume of CO2 :

At STP, molar volume = 22.4mol/L

moles of CO2 = "\frac{ volume}{molar volume}"

4.374mol = "\frac{volume}{22.4mol/L}"

Volume = 4.374mol ×22.4mol/L

Volume = 98mL

Volume of CO2 at STP, is 98mL